# Towards Generalizable Neural Solvers for Vehicle Routing Problems via Ensemble with Transferrable Local Policy

### v2

@@ 1 Introduction
Recently, machine learning has been adapted to solve combinatorial optimization problems Bengio et al. (2021); Cappart et al. (2023), which utilizes deep neural networks to extract features, learn heuristics, and generate solutions, named Neural Combinatorial Optimization (NCO) Bello et al. (2017). Depending on the training scheme, NCO includes Supervised Learning (SL) Joshi et al. (2019) and Reinforcement Learning (RL) methods Khalil et al. (2017); Kool et al. (2019). SL-based methods usually require numerous (nearly) optimal solutions as labels, which are hardly applicable if optimal solutions are expensive to obtain Bello et al. (2017). RL-based methods train the model using objective function evaluations as rewards, which are effective in most scenarios Zhang et al. (2020); Song et al. (2023).
@@
Towards generalizing to cross-distribution and cross-scale problem instances, several methods have been proposed recently. Some of them focus on generalization to large-scale problems. They often utilize the divide-and-conquer framework that can extend small-scale models to generalizable ones Fu et al. (2021); Hou et al. (2023); Cheng et al. (2023). Meanwhile, the inductive bias of symmetries Drakulic et al. (2023) and some novel neural architectures, such as heavy decoder Luo et al. (2023) and diffusion models Sun and Yang (2023), are found to be useful for cross-scale generalization. The other methods aim to address the cross-distribution issue. A simple solution is directly training the model on instances with various node distributions, which is effective but can be further improved by applying some advanced techniques, such as distributionally robust optimization Jiang et al. (2022), knowledge distillation Bi et al. (2022), graph contrastive learning Jiang et al. (2023), and meta-learning Manchanda et al. (2022). Note that the latest meta-learning method Zhou et al. (2023) considers an omni setting including both cross-distribution and cross-scale generalization.
@@
Considering that the general optimization objective of VRPs is to achieve the shortest routing length, the local neighborhood, containing adjacent nodes, plays an important role in the decision of moving from node to node. Meanwhile, local topological features have great potential to be transferable across both node distributions Jiang et al. (2023) and scales. To leverage the properties of local information, we propose a novel local policy that restricts the state and action space to a small number of local neighbor nodes. Furthermore, we integrate the local policy with a typical construction policy (e.g., POMO Kwon et al. (2020)) called global policy, which learns from the global information of complete VRP instances, to form an ensemble policy. The local and global policies are jointly trained to perform cooperatively and complementarily, achieving a good generalization.
@@ 2.3 Generalization Issue
Most existing methods are trained on small-scale problem instances (where $N\leq 100$) generated from uniform node distribution, and tested on instances from the same distribution. However, real-world problems are often with complex and unknown node distributions and large scales. This has led to the less satisfactory generalization performance of existing neural methods, which is recognized as a bottleneck for applications Joshi et al. (2022); Bi et al. (2022). Many related works have tried to alleviate the generalization issue by leveraging divide-and-conquer strategy Hou et al. (2023), meta learning Manchanda et al. (2022), symmetry of CO problems Drakulic et al. (2023), and heavy decoder architecture Luo et al. (2023), while most of them ignore either the cross-distribution or cross-scale setting.
@@
In this paper, we aim to address the generalization issue by ensemble with a transferrable local policy. Through training on uniform problem instances with problem scale $N=100$, we can achieve better generalization performance on both cross-distribution and cross-scale test instances, which will be clearly shown in our experiments. To effectively verify the generalization performance, we select the challenging CVRPLIB Uchoa et al. (2017); Arnold et al. (2019) and TSPLIB Reinelt (1991) as our test datasets since they cover diverse instances and many of them are from the real world, while previous synthetic test datasets with either fixed node distributions or limited problem scales can only cover a narrow subspace of the whole problem space.
@@ 3.1 Transferrable Local Policy
When solving VRPs, a common observation is that the optimal action (i.e., next node) is usually included in a small local neighborhood of the current node Helsgaun (2000), and the patterns of such local neighborhoods have the potential to be transferrable across various node distributions Jiang et al. (2023) and scales. Inspired by these facts, we propose to design a local MDP formulation that restricts the state and action space to the local neighborhood, aiming to achieve better generalization performance. The overall procedure of the proposed local policy is illustrated in Figure 2. Here we define the state and action space of local policy in details.
@@ 3.3 Ensemble of Global and Local Policies
where $N$ is the number of trajectories that is equal to the number of nodes, $B$ is the batch size, and $R_{i,j}$ is the reward of the $i$-th trajectory $\bm{\tau}_{i,j}$ on the $j$-th instance.
@@ 4 Experiments
We conduct experiments on two typical VRPs, i.e., Traveling Salesman Problem (TSP) and Capacitated VRP (CVRP). This section presents the experimental settings and results to evaluate the performance of our method and verify the efficacy of important components. Our code and data are available at https://github.com/gaocrr/ELG.
@@ 4.1 Experimental Settings
The training instances are randomly generated with uniform node distribution and the problem scale $N=100$, which is a common setting in previous works Nazari et al. (2018); Kwon et al. (2020). More details of training settings can be found in Appendix B.2. We consider the following test datasets to compare different methods.
@@ 4.2 Results on TSPLIB and CVRPLIB
Because the test distribution is complex and the problem scales of $N\leq 200$ are close to that of training instances, we can evaluate the cross-distribution performance by this subset. As shown in the second columns of Tables 1 and 2, ELG-POMO outperforms all the methods on the $N\leq 200$ subset, and the gap reductions on CVRP and TSP are $0.53\%$ and $0.50\%$, respectively, compared to the runner-up (Omni-POMO for CVRP, and BQ for TSP). These results clearly demonstrate the superiority of ELG-POMO in cross-distribution generalization, which will be further verified by the results in Table 4.
@@
The subset of $N\in(200,1000]$ (or $N\in(200,1002]$) contains instances with both cross-distribution and cross-scale characteristics. ELG-POMO reduces the gap from $6.95\%$ of Omni-POMO to $6.46\%$ for CVRP, and reduces the gap from $7.10\%$ of Omni-POMO to $5.60\%$ for TSP. Considering that the Omni-POMO is trained on cross-distribution (uniform and Gaussian mixture) and cross-scale ($N\in[50,200]$) instances while ELG-POMO only uses uniform instances with the problem scale $N=100$ in the training, the above comparison clearly demonstrates the better cross-distribution and cross-scale performance of the proposed method. Note that ELG-POMO outperforms BQ and LEHD on CVRP, but lags behind them on TSP with $N\in(200,1002]$. This may be attributed to their heavy decoder or decoder-only architectures, which can learn strong scale-invariant features Luo et al. (2023) to achieve good large-scale performance. We can also observe from Table 2 that ELG-POMO can surpass the diffusion mode-based methods, i.e., DIFUSCO and T2TCO, under the greedy inference setting.
@@
DIFUSCO 1.84% 10.83% 5.77% 1.73s
T2TCO 1.87% 9.72% 5.30% 1.71s
ELG-POMO (Ours) 1.12% 5.60% 3.08% 0.63s
@@ B.2 Implementation Details in Training
Following the common setting to generate training instances Nazari et al. [2018]; Kool et al. [2019], the node coordinates $\{(x_{i},y_{i})|\;i\in[N]\}$ are randomly generated from a uniform distribution $\mathcal{U}(0,1)$, the capacity $Q$ is set to $50$, and the demands $\{d_{i}|\;i\in[N]\}$ are randomly generated from a uniform distribution $\mathcal{U}(1,10)$.

### v3

@@ 1 Introduction
Recently, machine learning has been adapted to solve combinatorial optimization problems Bengio et al. (2021); Cappart et al. (2023), which utilizes deep neural networks to extract features, learn heuristics, and generate solutions, named Neural Combinatorial Optimization (NCO) Bello et al. (2017). Depending on the training scheme, NCO includes Supervised Learning (SL) Joshi et al. (2019); Chen et al. (2023) and Reinforcement Learning (RL) methods Kool et al. (2019). SL-based methods usually require numerous (nearly) optimal solutions as labels, which are hardly applicable if optimal solutions are expensive to obtain Bello et al. (2017). RL-based methods train the model using objective function evaluations as rewards, which are effective in most scenarios Zhang et al. (2020); Song et al. (2023).
@@
Towards generalizing to cross-distribution and cross-scale problem instances, several methods have been proposed recently. Some of them focus on generalization to large-scale problems. They often utilize the divide-and-conquer framework that can extend small-scale models to generalizable ones Fu et al. (2021); Hou et al. (2023); Cheng et al. (2023). Meanwhile, the inductive bias of symmetries Drakulic et al. (2023) and some novel neural architectures, such as heavy decoder Luo et al. (2023) and diffusion models Sun and Yang (2023), are found to be useful for cross-scale generalization. The other methods aim to address the cross-distribution issue. A simple solution is directly training the model on instances with various node distributions, which is effective but can be further improved by applying some advanced techniques, such as distributionally robust optimization Jiang et al. (2022), knowledge distillation Bi et al. (2022), graph contrastive learning Jiang et al. (2023b), and meta-learning Fu et al. (2023); Manchanda et al. (2022); Son et al. (2023). Note that the latest meta-learning method Zhou et al. (2023) considers an omni setting including both cross-distribution and cross-scale generalization.
@@
Considering that the general optimization objective of VRPs is to achieve the shortest routing length, the local neighborhood, containing adjacent nodes, plays an important role in the decision of moving from node to node. Meanwhile, local topological features have great potential to be transferable across both node distributions Jiang et al. (2023b) and scales. To leverage the properties of local information, we propose a novel local policy that restricts the state and action space to a small number of local neighbor nodes. Furthermore, we integrate the local policy with a typical construction policy (e.g., POMO Kwon et al. (2020)) called global policy, which learns from the global information of complete VRP instances, to form an ensemble policy. The local and global policies are jointly trained to perform cooperatively and complementarily, achieving a good generalization.
@@ 2.3 Generalization Issue
Most existing methods are trained on small-scale problem instances (where $N\leq 100$) generated from uniform node distribution, and tested on instances from the same distribution. However, real-world problems are often with complex and unknown node distributions and large scales. This has led to the less satisfactory generalization performance of existing neural methods, which is recognized as a bottleneck for applications Joshi et al. (2022); Bi et al. (2022); Liu et al. (2023). Many related works have tried to alleviate the generalization issue by leveraging divide-and-conquer strategy Hou et al. (2023), meta learning Manchanda et al. (2022), symmetry of CO problems Drakulic et al. (2023), and heavy decoder architecture Luo et al. (2023), while most of them ignore either the cross-distribution or scale setting.
@@
In this paper, we aim to address the generalization issue by ensemble with a transferrable local policy. Unlike a related ensemble method for NCO Jiang et al. (2023a) that employs the well-known Bootstrap method, we leverage different state and action spaces (i.e., local and global views) to generate diverse and complementary policies. To effectively verify the generalization performance, we select the challenging CVRPLIB Uchoa et al. (2017); Arnold et al. (2019) and TSPLIB Reinelt (1991) as our test datasets since they cover diverse instances and many of them are from the real world, while previous synthetic test datasets with either fixed node distributions or limited problem scales can only cover a narrow subspace of the whole problem space.
@@ 3.1 Transferrable Local Policy
When solving VRPs, a common observation is that the optimal action (i.e., next node) is usually included in a small local neighborhood of the current node Helsgaun (2000), and the patterns of such local neighborhoods have the potential to be transferrable across various node distributions Jiang et al. (2023b) and scales. Inspired by these facts, we propose to design a local MDP formulation that restricts the state and action space to the local neighborhood, aiming to achieve better generalization performance. The overall procedure of the proposed local policy is illustrated in Figure 2. Here we define the state and action space of local policy in details.
@@ 3.3 Ensemble of Global and Local Policies
where $N$ is the number of trajectories that is equal to the number of nodes, $B$ is the batch size, and $R_{i,j}$ is the reward of the $j$-th trajectory $\bm{\tau}_{i,j}$ on the $i$-th instance.
@@ 4 Experiments
We conduct experiments on two typical VRPs, i.e., Traveling Salesman Problem (TSP) and Capacitated VRP (CVRP). This section presents the experimental settings and results to evaluate the performance of our method and verify the efficacy of important components. Our code and data are available in https://github.com/lamda-bbo/ELG.
@@ 4.1 Experimental Settings
The training instances are randomly generated with uniform node distribution and the problem scale $N=100$, which is a common setting in previous works Kool et al. (2019); Kwon et al. (2020). More details of training settings can be found in Appendix B.2. We consider the following test datasets to compare different methods.
@@ 4.2 Results on TSPLIB and CVRPLIB
Because the test distribution is complex and the problem scales of $N\leq 200$ are close to that of training instances, we can evaluate the cross-distribution performance by benchmarks with $N\leq 200$. As shown in the second columns of Tables 1 and 2, ELG-POMO outperforms all the methods on the $N\leq 200$ subset, and the gap reductions on CVRP and TSP are $0.53\%$ and $0.50\%$, respectively, compared to the runner-up (Omni-POMO for CVRP, and BQ for TSP). These results clearly demonstrate the superiority of ELG-POMO in cross-distribution generalization, which will be further verified by the results in Table 4.
@@
The subset of $N\in(200,1000]$ (or $N\in(200,1002]$) contains instances with both cross-distribution and cross-scale characteristics. ELG-POMO reduces the gap from $6.95\%$ of Omni-POMO to $6.46\%$ for CVRP, and reduces the gap from $7.47\%$ of Omni-POMO to $5.90\%$ for TSP. Considering that the Omni-POMO is trained on cross-distribution (uniform and Gaussian mixture) and cross-scale ($N\in[50,200]$) instances while ELG-POMO only uses uniform instances with the problem scale $N=100$ in the training, the above comparison clearly demonstrates the better cross-distribution and cross-scale performance of the proposed method. Note that ELG-POMO outperforms BQ and LEHD on CVRP, but lags behind them on TSP with $N\in(200,1002]$. This may be attributed to their heavy decoder or decoder-only architectures, which can learn strong scale-invariant features Luo et al. (2023) to achieve good large-scale performance. We can also observe from Table 2 that ELG-POMO can surpass the diffusion mode-based methods, i.e., DIFUSCO and T2TCO, under the greedy inference setting.
@@
DIFUSCO 1.84% 10.83% 5.77% 30.68s
T2TCO 1.87% 9.72% 5.30% 30.82s
ELG-POMO (Ours) 1.12% 5.90% 3.08% 0.63s
@@ 5 Conclusion
Acknowledgments
The authors want to thank the anonymous reviewers for their helpful comments and suggestions. This work was supported by the National Science and Technology Major Project (2022ZD0116600) and National Science Foundation of China (62276124). Chao Qian is the corresponding author. The conference version of this paper has appeared at IJCAI’24.
@@ B.2 Implementation Details in Training
Following the common setting to generate training instances Kool et al. [2019]; Kwon et al. [2020], the node coordinates $\{(x_{i},y_{i})|\;i\in[N]\}$ are randomly generated from a uniform distribution $\mathcal{U}(0,1)$, the capacity $Q$ is set to $50$, and the demands $\{d_{i}|\;i\in[N]\}$ are randomly generated from a uniform distribution $\mathcal{U}(1,10)$.
@@ C.4 Comparison on CVRP Instances with Complex Distributions
C.5 Results of only local policy
The transferrable local policy is a critical component of our proposed method, which contributes greatly to both cross-distribution and cross-scale performance. Therefore, we evaluate the only local policy for clear ablation. The empirical results are shown in Table 12. For comparison fairness, we train a local policy with a similar number of parameters as the global policy by adding $3$ encoder layers and increasing the embedding dimension to $128$.
By comparing the results in the 2nd and 3rd rows, we observe that the local policy significantly improves the out-of-distribution performance on TSPLIB, verifying the good transferability of local policy. However, the local policy’s in-distribution performance on TSP100 deteriorates, probably because of the reduced state and action spaces.
As shown in Table 12, neither the local policy nor global policy performs worse than the proposed ELG-POMO method and these two policies exhibit complementary performances, validating the necessity of combining the local and global policies in our approach.
Methods TSP100 TSPLIB
Only local policy 1.71% 3.81%
Only global policy (POMO) 0.40% 7.45%
ELG-POMO 0.22% 3.08%
Table 12: Ablation results of only the local policy.